OpenStudy (anonymous):

r/12= 1 9/10 A. 22 4/5 B. 12 9/10 C. 10 4/5 D. 19/120

OpenStudy (anonymous):

what's this?

OpenStudy (anonymous):

r = 12 x19/ 10 = 6x19/5 =114/5

OpenStudy (anonymous):

i dony know